OpenStudy (anonymous):

12. (a) A series circuit has a capacitor of 1E−5 farad, a resistor of 3E2 ohms, and an inductor of 0.2 henry. The initial charge on the capacitor is 1E−6 coulomb and there is no initial current. Find the charge q on the capacitor at any time t. (b) Draw a phase portrait of the equivalent dynamical system that includes the trajectory corresponding to the initial value problem in part (a).

OpenStudy (anonymous):

\[Ldi/dt + Ri + (1/C)q = e(t)\]

OpenStudy (anonymous):

\[ i_{0}^{'}= \frac{e(t_0) - Ri_0 - (1/C)q_0 }{ L }\]

OpenStudy (anonymous):

\[\frac{ d ^{2} Q}{ dt ^{2} } + \frac{ R }{ L }\frac{ dQ}{ dt } + \frac{ 1 }{LC }Q=0\] \[\lambda ^{2} +\frac{ R }{ L }\lambda +\frac{ 1 }{ LC } = 0\]

OpenStudy (anonymous):

Forgot to say thanks, that helped.